OpenStudy (anonymous):

A cylindrical water tank has a hemispherical bottom. Find its total surface area and its volume if the dimensions are r=1.5 m h=3 m Please help me!

OpenStudy (lgbasallote):

\[\huge SA = 2\pi r^2 + 2\pi r h\] \[\huge V = \pi r^2 h\] does that help?

OpenStudy (lgbasallote):

where: SA = surface area r = radius h = height V = volume